# JWST spectroscopy of a blue binary cold classical Kuiper belt object

### v1

@@ 2.5 Reflectance Spectra
2001 XR254 was situated at the nominal pointing position near the center of the FOV during our observation, and as such, we expect the wavelength-dependent sensitivity profile of the extracted spectrum to closely mirror that of the SNAP-2 spectrum measured from the 20-point dither observation. On the other hand, the location of 2016 BP81 relative to the center of the FOV in all four dither positions is more similar to the region of the detector sampled by the 4-point nod observation of SNAP-2. We therefore paired the irradiance spectra of 2001 XR254 and 2016 BP81 with the 20-point dither and 4-point nod spectra of SNAP-2, respectively, when deriving the reflectance spectra.
@@
Lastly, we needed to account for the systematic difference between the spectra of SNAP-2 and the Sun. In Figure 2, we include the solar standard spectrum from the Planetary Spectrum Generator (PSG; G. L. Villanueva et al. 2018), which has been resampled and convolved to match the wavelength grid and spectral resolution of the NIRSpec PRISM grating. A marked discrepancy in spectral shape is seen at short wavelengths, reflecting a difference in temperature between the two stars. We corrected the reflectance spectra of our targets by multiplying the ratio between the spectra of SNAP-2 and the Sun. Figure 3 shows the final reflectance spectra of 2001 XR254 and 2016 BP81 (both components), normalized to unity at 2.5 µm.
2.6 Band Depth Calculations
@@
To facilitate quantitative comparisons between spectra, we calculated the depths of the main absorption features of H2O and CO2 ice, which lie at 2.0, 3.0, and 4.27 µm. For each band, we stipulated narrow wavelength ranges on one or both sides of the feature that we used to establish the continuum level across the band. We also established a central wavelength region in the middle of the band that we take as the band center. Next, we carried out an error-weighted linear least-squares fit to the points within the continuum region(s) to model the local continuum. This continuum model was then divided away from the reflectance spectrum.
@@
The band depth was computed by calculating the weighted average of all points within the band center region. The typical flux uncertainties provided by the empirical PSF fitting method in jwstspec (see Section 2.3) underestimate the true scatter of the data and cannot account for correlated noise across the wavelength axis. In order to derive more realistic band depth values and uncertainties, we calculated the ratio between the scatter in the residuals from the continuum fit and the median flux uncertainties within the continuum region(s). The flux uncertainty array of the spectrum was inflated by this ratio when calculating the band depth. The uncertainty in the band depth was set to the standard deviation of the data points within the band center region.
@@
For the 2-µm H2O-ice absorption, the continuum regions were set to 1.80–1.90 µm and 2.20–2.30 µm, and the band center range was taken to be 2.00-2.05 µm. Only the left side of the 3-µm H2O-ice band (2.60–2.65 µm) was used to establish a flat continuum level across the broad absorption; the band depth was calculated at 2.975–3.025 µm. The CO2-ice band depth at 4.26–4.28 µm was calculated using the continuum regions 4.00–4.15 and 4.35–4.45 µm.
@@ 3 Discussion
From an initial visual inspection of the reflectance spectra, we find that the primary and secondary components of 2016 BP81 have indistinguishable spectral profiles, with a positive continuum slope throughout the 0.7–2.5 µm wavelength range, a broad 3-µm H2O ice absorption band, hints of an additional weak H2O ice feature centered at 2 µm, a deep primary CO2-ice band at 4.27 µm, and clear evidence for the 2.7-µm CO2 combination band. The strong agreement between the binary components’ spectra indicates that the two bodies have the same surface composition. This finding is in line with models of binary formation that involve gravitational collapse, in which both components accrete out of a single self-gravitating pebble cloud (e.g., A. N. Youdin & J. Goodman, 2005; D. Nesvorný et al., 2010, 2019; A. H. Parker et al., 2011). Given the identical spectral profiles of the 2016 BP81 binary components, we averaged the two together and utilize the combined spectrum in the following discussion.
@@
In contrast, 2001 XR254 has a much steeper spectral slope below 1.0 µm, which levels off to a slightly negative-sloped continuum through 2.5 µm. At longer wavelengths, the right side of the broad 3-µm H2O-ice feature rises to significantly lower reflectance values than the left side, with a separate absorption band between 3.2 and 3.6 µm characteristic of aliphatic organics. In addition to deep CO2 features, 2001 XR254 also displays a strong CO ice absorption band centered at 4.7 µm. Minor features between 2.2 and 2.6 µm are suggestive of solid-phase CH3OH. Altogether, these distinct properties describe a surface composition very unlike that of the blue binary 2016 BP81 and is instead comparable to the CH3OH-rich compositional class of KBOs uncovered by N. Pinilla-Alonso et al. (2025). We therefore classify 2001 XR254 as a red binary cold classical object; indeed, the relatively imprecise measured visible spectral slope of %/100 nm reported in W. C. Fraser et al. (2021) straddles the threshold between the blue and red KBO subpopulations.
@@
To examine our two targets more closely within the broader context of known KBO surface types, we reprocessed the NIRSpec observations of KBOs obtained as part of the large Cycle 1 program presented in N. Pinilla-Alonso et al. (2025), following an identical methodology to the one we applied to our cold classical targets. We excluded all of the Centaurs (inward-scattered KBOs on orbits within the giant planet region), due to the likelihood of secondary evolution across their surfaces from the enhanced temperatures and space weathering they experience; strong evidence of this trend was recently demonstrated in a dedicated analysis of their JWST spectra (J. Licandro et al., 2025). Likewise, the exceptionally H2O-ice-rich Haumea collisional family member 2003 UZ117 was not considered. The size range spanned by this spectroscope sample is 100–800 km, which overlaps at the lower end with the diameters of the cold classical targets from our program. Figure 4 shows the results of this analysis, with all spectra normalized to unity at 2.5 µm and grouped by compositional class. The combined 2016 BP81 spectrum is overplotted in each panel with the black curve, while the spectrum of 2001 XR254 is displayed in purple alongside the other CH3OH-rich objects.
@@
From the figure, it is evident that the reflectance spectra of 2001 XR254 and 2016 BP81 generally lie within the range of spectral shapes attested across the broader CH3OH-rich and H2O-rich subclasses, respectively. The sample from N. Pinilla-Alonso et al. (2025) includes five other red cold classicals (three of them binaries), all of which belong to the CH3OH-rich compositional class, reinforcing our previous statement that 2001 XR254 is a red binary cold classical. Due to our stated focus on blue binaries in this paper, we do not consider 2001 XR254 in the following discussion. In the case of 2016 BP81, while its overall spectral profile is similar to that of the other H2O-rich KBOs, the distinctive combination of the relatively deep 4.27-µm CO2 band, the very weak 2-µm H2O-ice feature, and the absence of the 3.1-µm H2O-ice Fresnel peak appears to set it apart from the rest of its compositional class. Moreover, the reflectance level of the 2016 BP81 spectrum in the vicinity of 3.5 µm is substantially lower than the values spanned by the other spectra, suggesting that the surface of 2016 BP81 is enriched in aliphatic organics relative to the typical H2O-rich KBO.
@@
To investigate the relationship between 2016 BP81 and other H2O-rich KBOs in a more quantitative manner, we carried out a systematic band shape analysis across the H2O-rich compositional class. Here, we included some additional H2O-rich objects from other observing programs: (1) seven Neptune Trojans from Program #2550 (PI: L. Markwardt; L. Markwardt et al. 2023), and (2) three extreme transneptunian objects (ETNOs) from Program #4665 (PI: B. Holler). All of these datasets were passed through an analogous data reduction and spectral extraction workflow. We measured the band depths of the 2- and 3-µm H2O-ice absorptions, as well as the primary CO2 band at 4.27 µm, following the methodology described in Section 2.6.
@@
A grid plot of the two-parameter distributions is shown in Figure 5. Different symbols denote the respective dynamical subpopulations that the targets belong to, with the dynamical classifications taken from the recent work by K. Volk & C. Van Laerhoven (2024). Note that in some cases, the classifications here differ from those provided in N. Pinilla-Alonso et al. (2025).
The relationship between the 2- and 3-µm band depths is an important probe for studying variations in the relative abundance and physical properties of H2O ice across a sample of compositionally similar surfaces. Previous laboratory work (e.g., R. M. E. Mastrapa et al., 2013) and surface modeling of H2O-ice-rich solar system minor bodies (e.g., S. Protopapa et al., 2014; A. Raponi et al., 2016) have demonstrated that the two band depths vary in tandem when the overall H2O abundance changes; meanwhile, alterations to the size of H2O-ice grains manifest as changes to the relative depths between the two bands, with a higher 3- vs. 2-µm depth ratio indicative of smaller grain sizes. Looking at the top panel in Figure 5, we find that most of the H2O-rich KBOs trace out a positive correlation line, suggesting a range of relative H2O-ice abundances but otherwise similar particle size distributions.
@@
However, 2016 BP81 is visibly offset from this covariance trend and is situated within a tight cluster alongside the Neptune Trojan 2011 SO277 and the ETNO 2016 QV89 at higher 3- vs. 2-µm depth ratios, implying smaller H2O-ice grain sizes. The spectra of these three objects are plotted together in Figure 6, from which the similarity in H2O-ice band shapes is evident. The positions of the three objects in the band-depth grid plot are linked by dashed black lines to facilitate interpretation.
@@
When assessing the plots of the CO2 band depth as a function of the 2- and 3-µm H2O band depths in Figure 5, we once again observe a marked distinction between the blue binary 2016 BP81 and the rest of the H2O-rich KBOs. While most objects have relative CO2 band depths between 0.1 and 0.4 that are uncorrelated with their 2- or 3-µm H2O-ice absorption feature depths, both 2016 BP81 and 2011 SO277 have significantly deeper CO2-ice spectral signatures. Meanwhile, although 2016 QV89 does not stand out within the broader distribution of CO2 band depths, it nevertheless lies at the upper end of the range spanned by the other H2O-rich KBOs. From Figure 6, we can see how comparable the CO2 primary band shapes (and indeed the entire 3.5–5.1 µm spectral profiles) are among all three objects. Moreover, when examining the 3.2–3.6 µm wavelength range, we notice that 2016 QV89 exhibits a similar spectral shape to 2016 BP81, while 2011 SO277 displays an additional absorption band similar to what is seen on the CO2- and CH3OH-rich KBOs (see Figure 4). It follows that all three objects’ surfaces are enriched in aliphatic organics relative to the rest of the H2O-rich KBOs.
@@
The discrepant spectroscopic behavior of 2016 BP81, 2011 SO277, and 2016 QV89 points toward a distinct surface subtype within the wider H2O-rich compositional class that is characterized by systematically finer H2O-ice grains, more abundant CO2 ice, and stronger signatures of aliphatic organics. We consider differences in current thermal and space weathering environment to be an unlikely explanation for this distinction, given that the three objects have drastically different present-day orbits. 2011 SO277 is particularly diagnostic in this regard, as it lies within the very narrow range of average surface temperatures and energetic particle fluences spanned by the 1:1 mean motion resonance with Neptune, while being compositionally distinct from the other six H2O-rich Neptune Trojans in our sample. Instead, we posit that the distinctive surface properties may reflect differences in the objects’ initial formation environment.
@@
With only one blue binary in the combined spectroscopic sample, it is impossible to make general statements about the blue binary cold classical population. Observations of additional targets are necessary to determine if the distinctive properties of 2016 BP81 are a characteristic trait of the population, or if instead the blue binaries include objects with surface properties that are more typical of the broader H2O-rich KBO subclass (as is the case with the Neptune Trojans and ETNOs). The answer to that question has fundamental implications for our understanding of the origin and dynamical evolution of the blue binaries.
@@
D. Nesvorný et al. (2022) put forward two proposals to explain the presence of blue binaries within the cold classical region (see Introduction), and we now revisit them with the observed properties of 2016 BP81 in mind. In the gentle push-out scenario, dynamical modeling has demonstrated that the survival of the most widely-separated blue binaries requires an initial source region beyond 30 au, which necessarily places the primordial blue-to-red color transition at even larger heliocentric distances. Meanwhile, the overall implantation efficiency increases with increasing initial orbital semimajor axis, which means that the current blue binary population should preferentially sample the region of the planetesimal disk nearest to the color transition. From the spectrum of 2016 BP81, we do find distinctive spectral features that could plausibly be interpreted as being transitional between those typical of the blue H2O-rich KBOs and the redder CO2- and CH3OH-rich compositional classes. These include the exceptionally strong CO2 primary band and the higher abundance of aliphatic organics, as well as the relatively weak 2-µm H2O-ice absorption (see Figure 4).
@@
The push-out hypothesis allows for cold classical implantation from a wide range of initial heliocentric distances within the outer protoplanetary disk (e.g., D. Nesvorný et al., 2020). Therefore, one may expect future observations of blue binaries to reveal a range of surface compositions, including those that are more consistent with the majority of H2O-rich KBOs in our spectroscopic sample. Within this framework, 2011 SO277 and 2016 QV89 would also be interpreted as originating from the blue-to-red transition region prior to emplacement into their respective present-day dynamical populations. However, we recall that the simulations of the push-out scenario presented in D. Nesvorný et al. (2022) predict significant implantation of blue singletons, which is inconsistent with our current knowledge of the blue cold classical population.
@@
Meanwhile, our spectroscopic sample presents another challenge to the push-out scenario—the apparent absence of CO2-rich objects within the cold classical population. Under the compositional gradient framework described in N. Pinilla-Alonso et al. (2025), which places the initial formation region of the CO2-rich KBOs in between that of the H2O- and CH3OH-rich subclasses, one would expect the outward implantation of the H2O-rich blue binaries from the inner transneptunian disk to be accompanied by a significant admixture of CO2-rich objects in the red cold classical population. However, all six red cold classicals with published JWST spectra (including 2001 XR254 from our observations) belong to the CH3OH-rich compositional class. In fact, out of all dynamical subpopulations within the Kuiper belt, only the cold classicals appear to lack CO2-rich objects (M. N. De Prá et al., 2025; N. Pinilla-Alonso et al., 2025). Without future refinements to the dynamical models that address the paucity of blue singletons and CO2-rich objects within the cold classical population, the push-out hypothesis appears to be untenable.
@@
The alternative scenario described in D. Nesvorný et al. (2022) is the in situ formation of blue binaries at 42–47 au during the early stages of the protoplanetary disk, when temperatures in the disk midplane were higher and fewer volatile ices (e.g., methane and carbon monoxide) were in the solid phase (e.g., B. Bitsch et al., 2015). Objects that formed later, after the dispersion of the gaseous component of the disk and subsequent dissipative cooling of the disk midplane, would have accreted more volatile ices onto their surfaces and thereby developed the redder, CH3OH-rich surfaces typical of most cold classicals. Crucially, this proposed evolutionary pathway obviates the need to invoke dynamical implantation of the blue binaries into the cold classical region, and as such, the absence of blue singletons does not present a problem for the model. Moreover, the higher gas-to-dust ratio of the young protoplanetary disk would have facilitated the types of aerodynamic interactions that are conducive to binary formation via streaming instabilities and subsequent gravitational collapse (A. N. Youdin & J. Goodman, 2005; J. Squire & P. F. Hopkins, 2018; D. Nesvorný et al., 2021). It follows that the in situ formation scenario may be the key to explaining the near 100% binary fraction of the blue cold classicals (W. C. Fraser et al., 2017, 2021).
@@
From a compositional perspective, the unique conditions in the outermost reaches of the young protoplanetary disk could explain the distinctive spectral features of 2016 BP81 when compared to other H2O-rich KBOs, which may have instead formed at a later time and at smaller heliocentric distances—within the warmer inner region of the transneptunian planetesimal disk. If this scenario is correct, one would expect all blue binary cold classicals to display the same characteristic compositional signatures as 2016 BP81. Consequently, the spectrally similar objects 2011 SO277 and 2016 QV89 would likely have initially formed alongside the blue binaries before being scattered from the cold classical region, perhaps due to dynamical stirring of the outer protoplanetary disk during the final phases of Neptune’s outward migration.

### v2

@@ 2.5 Reflectance Spectra
During our observation, 2001 XR254 was situated at the nominal pointing position near the center of the FOV, and as such, we expect the wavelength-dependent sensitivity profile of the extracted spectrum to closely mirror that of the SNAP-2 spectrum measured from the 20-point dither observation. On the other hand, the location of 2016 BP81 relative to the center of the FOV in all four dither positions is more similar to the region of the detector sampled by the 4-point nod observation of SNAP-2. We therefore paired the irradiance spectra of 2001 XR254 and 2016 BP81 with the 20-point dither and 4-point nod spectra of SNAP-2, respectively, when deriving the reflectance spectra.
@@
Lastly, we needed to account for the systematic difference between the spectra of SNAP-2 and the Sun. In Figure 2, we include the solar standard spectrum from the Planetary Spectrum Generator (PSG; psg), which has been resampled and convolved to match the wavelength grid and spectral resolution of the NIRSpec PRISM grating. A marked discrepancy in spectral shape is seen at short wavelengths, reflecting a difference in temperature between the two stars. We corrected the reflectance spectra of our targets by multiplying the ratio between the spectra of SNAP-2 and the Sun. Figure 3 shows the final reflectance spectra of 2001 XR254 and 2016 BP81 (both components), normalized to unity at 2.5 m.
2.6 Band-depth Calculations
@@
To facilitate quantitative comparisons between spectra, we calculated the depths of the main absorption features of H2O and CO2 ice, which lie at 2.0, 3.0, and 4.27 m. For each band, we stipulated narrow wavelength ranges on one or both sides of the feature that we used to establish the continuum level across the band. We also established a central wavelength region in the middle of the band that we take as the band center. Next, we carried out an error-weighted linear least-squares fit to the points within the continuum region(s) to model the local continuum. This continuum model was then divided away from the reflectance spectrum.
@@
The band depth was computed by subtracting the weighted average of all points within the band center region from one. The typical flux uncertainties provided by the empirical PSF fitting method in jwstspec (see Section 2.3) underestimate the true scatter of the data and cannot account for correlated noise across the wavelength axis. In order to derive more realistic band-depth values and uncertainties, we calculated the ratio between the scatter in the residuals from the continuum fit and the median flux uncertainties within the continuum region(s). The flux uncertainty array of the spectrum was inflated by this ratio when calculating the band depth. The uncertainty in the band depth was set to the standard deviation of the data points within the band center region.
@@
For the 2 m H2O-ice absorption, the continuum regions were set to 1.80–1.90 and 2.20–2.30 m, and the band center range was taken to be 2.00–2.05 m. Only the left side of the 3 m H2O-ice band (2.60–2.65 m) was used to establish a constant continuum level across the broad absorption; the band depth was calculated at 2.975–3.025 m. The CO2-ice band depth at 4.26–4.28 m was calculated using the continuum regions 4.00–4.15 and 4.35–4.45 m.
@@ 3 Discussion
From an initial visual inspection of the reflectance spectra, we find that the primary and secondary components of 2016 BP81 have indistinguishable spectral profiles, with a positive continuum slope throughout the 0.7–2.5 m wavelength range, a broad 3 m H2O-ice absorption band, an additional weak H2O-ice feature centered at 2 m, a deep primary CO2-ice band at 4.27 m, and clear evidence for the 2.7 m CO2 combination band. The strong agreement between the binary components’ spectra indicates that the two bodies have the same surface composition. This finding is in line with models of binary formation that involve gravitational collapse, in which both components accrete out of a single self-gravitating pebble cloud (e.g., youdin2005; nesvorny2010; nesvorny2019b; parker2011). Given the identical spectral profiles of the 2016 BP81 binary components, we averaged the two together and utilize the combined spectrum in the following discussion.
@@
Meanwhile, the spectrum 2001 XR254 strongly contrasts with that of 2016 BP81. It is characterized by a much steeper spectral slope below 1.0 m, which transitions sharply to a negative-sloped continuum through 2.5 m. At longer wavelengths, the right side of the broad 3 m H2O-ice feature rises to significantly lower reflectance values than the left side, with a distinct absorption band between 3.3 and 3.6 m characteristic of aliphatic organics. In addition to deep CO2 features, 2001 XR254 also displays a strong CO-ice absorption band centered at 4.7 m, which is absent on 2016 BP81. Minor features between 2.2 and 2.6 m are suggestive of solid-phase CH3OH. Altogether, these distinct properties describe a surface composition very unlike that of the blue binary 2016 BP81 and are instead fully consistent with the organics-type compositional class of KBOs uncovered by pinillaalonso2025, all the members of which belong to the red color subpopulation. We therefore classify 2001 XR254 as a red binary cold classical object; indeed, the relatively imprecise measured visible spectral slope of %/100 nm reported in fraser2021 straddles the threshold between the less-red and red KBO subpopulations.
@@
To examine our two targets more closely within the broader context of known KBO surface types, we reprocessed the NIRSpec observations of KBOs obtained as part of the large Cycle 1 program presented in pinillaalonso2025, following an identical methodology to the one we applied to our cold classical targets. We excluded all of the Centaurs (inward-scattered KBOs on orbits within the giant planet region), due to the likelihood of secondary evolution across their surfaces from the enhanced temperatures and space weathering they experience; strong evidence of this trend was recently demonstrated in a dedicated analysis of their JWST spectra (licandro2025). Likewise, the exceptionally H2O-ice-rich Haumea collisional family member 2003 UZ117 was not considered. The size range spanned by this spectroscope sample is 100–800 km, which overlaps at the lower end with the diameters of the cold classical targets from our program. Figure 4 shows the results of this analysis, with all spectra normalized to unity at 2.5 m and grouped by compositional class. The combined 2016 BP81 spectrum is overplotted in each panel with the black curve, while the spectrum of 2001 XR254 is displayed in purple alongside the other organics-type objects.
@@
From the figure, it is evident that the reflectance spectra of 2001 XR254 and 2016 BP81 are largely consistent with the range of spectral shapes attested across the organics- and H2O-type compositional classes, respectively. In the case of 2001 XR254, all of its spectral properties—the steep spectral slope below 1.0 m, the negative-sloped continuum between 1.2 and 2.5 m, the presence of CH3OH ice features, the overall shape of the broad 3 m absorption, the clear presence of aliphatic organics, the shape of the CO2-ice band, and the profile of the CO-ice feature at 4.7 m—are archetypal of the organics-type KBOs. In particular, the contribution of CH3OH in the surface composition is a property that, while not shared by all organics-type KBOs, is only attested among members of that compositional class (brunetto2025). We also note that the sample from pinillaalonso2025 includes five other red cold classicals (three of them binaries), all of which are organics-type KBOs, reinforcing our previous statement that 2001 XR254 is a red binary cold classical. Due to our stated focus on blue binaries in this paper, we do not consider 2001 XR254 in the following discussion.
For 2016 BP81, the overall continuum shape, absence of CO ice, and weak CO2 combination band in its spectrum are broadly consistent with the H2O-type KBO class. However, several aspects of its spectrum suggest that the blue binary may represent a distinct type of surface composition that, while generally similar to the H2O-type objects shown in Figure 4, is somewhat adjacent in the space of spectral properties. It displays a relatively deep 4.27 m CO2 band, a very weak 2 m H2O-ice feature, and lacks the 3.1 m H2O-ice Fresnel peak. While each of these characteristics is present on other H2O-type KBOs in the sample from pinillaalonso2025, the combination of all three appears to be unique. Moreover, the reflectance level of the 2016 BP81 spectrum in the vicinity of 3.5 m is substantially lower than the values spanned by the other spectra, suggesting that the surface of 2016 BP81 may be enriched in aliphatic organics relative to the typical H2O-type KBO and/or characterized by a different ratio of amorphous vs. crystalline H2O ice (see the discussion below).
@@
To investigate the relationship between 2016 BP81 and other H2O-type KBOs in a more quantitative manner, we carried out a systematic band shape analysis across the H2O-type subpopulation. Here, we included some additional H2O-type objects from other observing programs: (1) seven Neptune Trojans from Program #2550 (PI: L. Markwardt; markwardt2023), and (2) three extreme trans-Neptunian objects (ETNOs) from Program #4665 (PI: B. Holler). All of these datasets were passed through an analogous data reduction and spectral extraction workflow. We measured the band depths of the 2 and 3 m H2O-ice absorptions, as well as the primary CO2 band at 4.27 m, following the methodology described in Section 2.6.
@@
The grid of two-parameter distribution plots is shown in Figure 5. Different symbols denote the respective dynamical subpopulations that the targets belong to, with the dynamical classifications taken from the recent work by volk2024.
@@
The relationship between the 2 and 3 m band depths is an important probe for studying variations in the relative abundance and physical properties of H2O ice across a sample of compositionally similar surfaces. Looking at the top panel in Figure 5, we find that the 2 and 3 m H2O-ice band depths for most of the H2O-type KBOs trace out a positive correlation line. However, 2016 BP81 is visibly offset from this trend and is situated within a tight cluster alongside the Neptune Trojan 2011 SO277 and the ETNO 2016 QV89 at higher 3 vs. 2 m depth ratios. The spectra of these three objects are plotted together in Figure 6, from which the similarity in H2O-ice band shapes is evident. The positions of the three objects in the band-depth grid plot are linked by dashed black lines to facilitate interpretation.
Previous laboratory work (e.g., mastrapa2009; mastrapa2013) and surface modeling of H2O-ice-rich solar system minor bodies (e.g., protopapa2014; raponi2016) have demonstrated that the 2 and 3 m band depths scale in tandem when the relative H2O-ice abundance on the surface is varied. Meanwhile, changes to the H2O-ice grain size modulate the relative depths between the two bands, with a higher 3 vs. 2 m depth ratio indicative of smaller grain sizes. While at face value the systematically enhanced 3 vs. 2 m band-depth ratios of 2016 BP81, 2011 SO277, and 2016 QV89 suggest surfaces with smaller H2O-ice grain sizes than on the other H2O-type KBOs, other possible confounding factors warrant caution when considering this interpretation.
Based on laboratory measurements of the wavelength-dependent absorption coefficient of H2O ice, the 2 m band probes significantly deeper penetration depths than the 3 m feature (e.g., mastrapa2013). It follows that vertical stratification within the uppermost layers of the surface can dramatically alter the relative depths of the H2O-ice bands, independent of the size of the ice grains (see, for example, the discussion in protopapa2024 related to compositional modeling of the JWST/NIRSpec spectrum of Pluto’s satellite Charon). Similarly, the specific configuration in which the H2O ice is incorporated into the surface regolith—whether as spatially segregated units or within intimate mixtures with other ices and/or opaque materials—has a major impact on the manner in which the various molecular features are expressed in the reflectance spectrum.
Another key physical property of H2O ice is crystallinity. While the mixing ratio between the crystalline and amorphous phases of H2O ice does not have as large of an impact as grain size on the relative band depths, it does fundamentally set the shape of the 2 and 3 m features (e.g., mastrapa2013; protopapa2014). In particular, the presence of a prominent Fresnel reflectance peak, as seen on many of the H2O-type KBOs plotted in Figure 4, is usually interpreted as an indicator of crystalline H2O ice (e.g., brown2006; pinillaalonso2025). It follows that the absence of such a peak in the spectrum of 2016 BP81 (as well as other H2O-type objects) suggests a surface where amorphous H2O ice is more dominant. However, given that a sharp Fresnel peak centered at 3.1 m requires grains of pure crystalline water ice that are larger than a few microns (e.g., mastrapa2009; stephan2021), the strength of this feature may be attenuated due to impurities in the ice grains (e.g., mixtures with other compounds such as CO2), very small grain sizes, or vertical stratification. Detailed spectral modeling, which is beyond the scope of this paper, is necessary to robustly disentangle the effects of grain size, crystallinity, mixing, and layering on the H2O-ice features in these spectra.
When examining the 3.3–3.6 m wavelength range, we find that the normalized spectrum of 2016 BP81 lies at a much lower reflectance level than the average H2O-type KBO. As illustrated in Figure 6, 2016 QV89 exhibits a similar spectral shape to 2016 BP81 in this region, while 2011 SO277 displays an additional characteristic inflection near 3.3 m that signals the presence of a distinct absorption band due to aliphatic organics, similar to the analogous features seen on CO2- and organics-type KBOs (see Figure 4). Given that the spectra of these three objects show a tight affinity throughout the rest of the NIRSpec wavelength range, the confirmed signature of organics on 2011 SO277 suggests that the surfaces of all three objects may be systematically enriched in aliphatic organics relative to the other H2O-type KBOs. Thus, when specifically considering the abundance of organics, these objects could be interpreted as being transitional between the typical H2O-type KBOs and the other two compositional classes. However, the distinct spectral profile of these three objects in this region could also be at least partially explained by a difference in H2O-ice crystallinity. Specifically, when controlling for grain size, the curvature and maximum reflectance level of the H2O-ice model spectra in the 3.3–3.6 m range are highly sensitive to crystallinity, with the crystalline phase showing systematically lower reflectance levels than the amorphous phase (mastrapa2009; protopapa2014). In light of this ambiguity in the interpretation of the spectra, we consider the relative enhancement of aliphatic organics on the blue binary 2016 BP81 to be tentative.
@@
When assessing the plots of the CO2 band depth as a function of the 2 and 3 m H2O band depths in Figure 5, we once again observe a marked distinction between the blue binary 2016 BP81 and the rest of the H2O-type KBOs. While most objects have relative CO2 band depths between 0.1 and 0.4 that are uncorrelated with their 2- or 3-m H2O-ice absorption feature depths, both 2016 BP81 and 2011 SO277 have significantly deeper CO2-ice spectral signatures. Meanwhile, although 2016 QV89 does not stand out within the broader distribution of CO2 band depths, it nevertheless lies at the upper end of the range spanned by the other H2O-type KBOs. From Figure 6, we can see how comparable the CO2 primary band shapes (and indeed the entire 3.5–5.1 m spectral profiles) are among all three objects.
@@
The discrepant spectroscopic behavior of 2016 BP81, 2011 SO277, and 2016 QV89 points toward a distinct surface subtype that is adjacent to the range spanned by typical H2O-type KBOs within the space of spectral properties. We consider differences in current thermal and space weathering environment to be an unlikely explanation for this distinction, given that the three objects have drastically different present-day orbits. Here, 2011 SO277 is particularly diagnostic in this regard, as it lies within the very narrow range of average surface temperatures and energetic particle fluences spanned by the 1:1 mean motion resonance with Neptune while being compositionally distinct from the other six H2O-type Neptune Trojans in our sample. Instead, we posit that the distinctive surface properties may reflect differences in the objects’ initial formation environment.
@@
With only one blue binary in the combined spectroscopic sample, it is impossible to make general statements about the blue binary cold classical population. Observations of additional targets are necessary to determine if the distinctive properties of 2016 BP81 are a characteristic trait of the population or if instead the blue binaries include objects with surface properties that are more typical of the broader H2O-type KBO class (as is the case with the Neptune Trojans and ETNOs). The answer to that question has fundamental implications for our understanding of the origin and dynamical evolution of the blue binaries.
@@
nesvorny2022 put forward two proposals to explain the presence of blue binaries within the cold classical region (see the Introduction), and we now revisit them with the observed properties of 2016 BP81 in mind. In the gentle pushout scenario, dynamical modeling has demonstrated that the survival of the most widely separated blue binaries requires an initial source region beyond 30 au, which necessarily places the primordial less-red/red color transition at even larger heliocentric distances. Meanwhile, the overall implantation efficiency increases with increasing initial orbital semimajor axis, which means that the current blue binary population should preferentially sample the region of the planetesimal disk nearest to the color transition. From the spectrum of 2016 BP81, we do find distinctive spectral features that could plausibly be interpreted as being transitional between those typical of the less-red H2O-type KBOs and the redder CO2- and organics-type compositional classes. These include the exceptionally strong CO2 primary band and the potentially higher abundance of aliphatic organics, as well as the relatively weak 2 m H2O-ice absorption (see Figure 4).
@@
The pushout hypothesis allows for cold classical implantation from a wide range of initial heliocentric distances within the outer protoplanetary disk (e.g., nesvorny2020). Therefore, one may expect future observations of blue binaries to reveal a range of surface compositions, including those that are more consistent with the majority of H2O-type KBOs in our spectroscopic sample. Within this framework, 2011 SO277 and 2016 QV89 would also be interpreted as originating from the less-red/red transition region prior to emplacement into their respective present-day dynamical populations. However, we recall that the simulations of the pushout scenario presented in nesvorny2022 predict significant implantation of less-red singletons, which is inconsistent with our current knowledge of the cold classical population.
@@
Meanwhile, our spectroscopic sample presents another challenge to the pushout scenario—the apparent absence of CO2-type objects within the cold classical population. Under the compositional gradient framework described in pinillaalonso2025, which places the initial formation region of the CO2-type KBOs in between that of the H2O- and organics-type classes, one would expect the outward implantation of the H2O-type blue binaries from the inner trans-Neptunian disk to be accompanied by a significant admixture of CO2-type objects in the red cold classical population. However, all six red cold classicals with published JWST spectra (including 2001 XR254 from our observations) belong to the organics-type compositional class. In fact, out of all dynamical subpopulations within the Kuiper belt, only the cold classicals appear to lack CO2-type objects (depra2025; pinillaalonso2025). Without future refinements to the dynamical models that address the paucity of less-red singletons and CO2-type objects within the cold classical population, the pushout hypothesis appears to be untenable.
@@
The alternative scenario described in nesvorny2022 is the in situ formation of blue binaries at 42–47 au during the early stages of the protoplanetary disk, when temperatures in the disk midplane were higher and fewer volatile ices (e.g., methane and carbon monoxide) were in the solid phase (e.g., bitsch2015). Objects that formed later, after the dispersion of the gaseous component of the disk and subsequent dissipative cooling of the disk midplane, would have accreted more volatile ices onto their surfaces and thereby developed the redder, organics-type surfaces typical of most cold classicals. Crucially, this proposed evolutionary pathway obviates the need to invoke dynamical implantation of the blue binaries into the cold classical region, and as such, the absence of less-red singletons does not present a problem for the model. Moreover, the higher gas-to-dust ratio of the young protoplanetary disk would have facilitated the types of aerodynamic interactions that are conducive to binary formation via streaming instabilities and subsequent gravitational collapse (youdin2005; squire2018; nesvorny2021). It follows that the in situ formation scenario may be the key to explaining the near-100% binary fraction of the less-red cold classicals (fraser2017; fraser2021).
@@
From a compositional perspective, the unique conditions in the outermost reaches of the young protoplanetary disk could explain the distinctive spectral features of 2016 BP81 when compared to other H2O-type KBOs, which may have instead formed at a later time and at smaller heliocentric distances—within the warmer inner region of the trans-Neptunian planetesimal disk. If this scenario is correct, one would expect all blue binary cold classicals to display the same characteristic compositional signatures as 2016 BP81. Consequently, the spectrally similar objects 2011 SO277 and 2016 QV89 would likely have initially formed alongside the blue binaries before being scattered from the cold classical region, perhaps due to dynamical stirring of the outer protoplanetary disk during the final phases of Neptune’s outward migration.